# Operation of silicon photomultipliers in a dilution refrigerator down to 9.4 mK towards a cryogenic cosmic ray muon veto system

### v1

@@ 3 Characterisation analysis
In this section, we describe the characterisation measurements performed on waveform data acquired from the NUV-HD-cryo SiPM whilst operated at $T_{\mathrm{MC}}$.
@@ 3.1 Single photon response and gain
We first characterise the single photoelectron (PE) response of the SiPM. To do this, we plot two histograms corresponding to the prompt pulse charge and pulse amplitude respectively of all found pulses in a given run. Figure 5 shows the amplitude (left) and prompt charge (right) distributions of the SiPM whilst operated at $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 4.8 V). In both histograms, there are distinct peaks for successive numbers of PE. The first peak is the 1 PE signal, and corresponds to the integrated charge or amplitude of the observed output current pulse following the detection of a single photon. The pedestal peak is not visible in our distributions since a requirement is made on the pulse height in the pulse finding algorithm applied to the data, to limit identifying false pulses originating from electronic noise.
@@
Due to intrinsic noise, a SiPM still generates output pulses in the absence of light. The 1 PE peak we observe in our distributions is due to a combination of thermally-induced noise called “dark counts” and noise from “correlated delayed avalanches”. These noise sources are discussed in detail in Sections 3.2 and 3.4 respectively. The successive peaks corresponding to >>1 PE we observe in our distributions are a result of correlated noise from “direct crosstalk”, which manifest as superpositions of the 1 PE waveform. This noise source is discussed further in Section 3.3.
@@
The single PE response is determined by fitting a gaussian to the first peak in each spectrum and extracting the fitted mean. The amplitude and prompt charge of the single PE peak at $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 4.8 V) is calculated to be $\bar{A}_{\mathrm{1PE}}$ = 3.65 $\pm$ 0.42 mV and $\bar{Q}_{\mathrm{1PE}}$ = 0.15 $\pm$ 0.02 Vs respectively. The single PE amplitude is smaller than the trigger level used during acquisition (5 mV) due to the moving average filter which is applied to the data prior to running the pulse finding algorithm, which reduces the height of the pulses; this can be seen in Figure 4.
@@
Figure 5: Pulse amplitude (left) and pulse prompt charge (right) distributions of the NUV-HD-cryo SiPM operated at $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 4.8 V). Gaussian functions are fit to the first 3 peaks in each histogram in order to extract the gain measured in amplitude (left) and charge (right). The first peak in each histogram corresponds to the 1 PE peak.
@@
With the single PE response, we can now calculate the gain, defined as the amount of charge generated in an avalanche in response to each detected photon. Each microcell in the SiPM produces a highly quantised amount of charge when a photon is absorbed in the active volume, and this response is extremely uniform across all microcells. This unique property of SiPMs results in excellent charge resolution, and allows for a very precise measurement of the gain. The gain is expected to be proportional with $\Delta V$ [hamamatsu].
The separation between each pair of consecutive peaks in Figure 5 is constant and constitutes the gain. To determine the separation, we fit individual gaussian functions to the remaining peaks in each spectrum. We compute the gain as the difference in the fitted means for the first and second peaks, $G=\mu_{2\mathrm{PE}}-\mu_{1\mathrm{PE}}$. The gain measured from the amplitude and charge distributions at $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 4.8 V) are $G_{A}$ = 3.52 $\pm$ 0.006 mV and $G_{Q}$ = 0.149 $\pm$ 0.0003 Vs respectively.
Figure 6 shows the measured gain as a function of $V_{\mathrm{bias}}$ ($\Delta V$) calculated using charge (blue) and amplitude (red) respectively. The gain calculated from the charge, $G_{Q}$, can be converted to the gain in units of the number of charge carriers produced in an avalanche, $G_{N_{e}}$, using the following equation:
@@
$G_{N_{e}}=\frac{G_{Q}\cdot t_{S}}{R_{\mathrm{scope}}\cdot G_{\mathrm{amp}}\cdot q_{e}},\qquad$ (3.1)
@@
Figure 6: Gain of NUV-HD-cryo SiPM in charge in $10^{5}N_{e}$] (blue), and amplitude in [mV] (red) as a function of $V_{\mathrm{bias}}$ ($\Delta V$). Solid red and dashed blue lines are interpolated between the data points. Error bars are shown, however are not necessarily visible as the errors are small.
3.2 Dark count rate
SiPMs can exhibit both uncorrelated and correlated noise contributions. The dark count rate (DCR) is the main source of uncorrelated noise. Each dark count is initiated by a thermally-generated charge carrier, which triggers an avalanche in the active region through the same process as a photon-generated charge carrier; thus DCR events are indistinguishable from single photon signals. The DCR of the SiPM is expected to increase with $\Delta V$ [NUV-HD-cryo], since a higher $\Delta V$ increases the probability of an avalanche being triggered. The DCR is also expected to increase as a function of temperature. At higher temperatures, the DCR is dominated by field-enhanced thermal generation of charge carriers. However, below a certain temperature, the reduction of the DCR reaches a plateau, where the DCR becomes dominated by tunnelling effects that are weakly temperature dependent.
We characterise the DCR of the NUV-HD-cryo SiPM using the unshadowing method, based on using the time distribution between two consecutive pulses described in [UnshadowingMethod]. Given the incredibly low DCR of this technology, we also consider the time differences of pulses across separate waveforms, assuming that a negligible number of pulses are lost between waveforms. This assumption can be made due to the data being acquired using the Tetronix FastFrame Memory scheme [FastFrame] described earlier in Section 2.3, which has a maximum dead time of 2.5 µs.
@@ 3.2 Dark count rate
To select primary pulses for this analysis, we first apply the pulse finding algorithm to the data to identify all candidate $\geq$1 PE pulses. We also require that the time difference between the primary pulse and the previous pulse is $\geq$ 50 µs. This cut is made to reduce the probability of the secondary pulse not being associated with the selected primary pulse to be negligible. Figure 7 (left) shows the measured time distribution of consecutive pulses whilst the SiPM is operated at $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 4.8 V). The dip in the $dt$ distribution at close to the waveform acquisition window, 100 µs, can be primarily attributed to a loss in the pulse finder efficiency for pulses that are spread across two continuous waveforms.
@@
Figure 7 (right) shows the unshadowed time distribution following the procedure in [UnshadowingMethod]. The unshadowed distribution is the average number of secondary pulses occurring between the bin limits of $t_{i}$ and $t_{i+1}$ defined in Figure 7 (left), normalised per unit time to determine the pulse rate. The pulse rate distribution contains two contributions: the correlated delayed pulse rate, and the DCR. At large times, the rate of correlated delayed pulses is expected to disappear, leaving only the DCR. To extract the DCR, we compute the weighted mean of the asymptotic rate, between 1 s <dt<<dt< 8 s. We calculate the DCR to be 3.89 $\pm$ 0.19 mHz/mm2 at $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 4.8 V), the value of which is indicated by the horizontal dashed line in Figure 7 to guide the eye.
Figure 8 shows the measured DCR as a function of $V_{\mathrm{bias}}$ ($\Delta V$). As expected, the DCR increases as a function of $\Delta V$, increasing by 25% for 4.8 V $\Delta V<$ 6.8 V. These values are also consistent with the measured DCR value of <<5 mHz/mm2 at 77 K from [NUV-HD-cryo] and are compatible with the observation that for temperatures below 77 K, the DCR is dominated by tunnelling effects which are weakly temperature-dependent and thus and in this regime, the DCR is approximately constant for falling temperature.
@@
Figure 7: Left: Measured time distribution (dtdt) of consecutive pulses at $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 4.8 V). Right: Unshadowed time distribution, based on dtdt distribution in left-hand plot. Horizontal dashed line indicates calculated DCR of 3.89 $\pm$ 0.19 mHz/mm2.
@@
Figure 8: Measured DCR of NUV-HD-cryo SiPM as a function of $V_{\mathrm{bias}}$ ($\Delta V$). The DCR increases by 25% between 4.8 V $\Delta V<$ 6.8 V. Red line is interpolated between the data points.

### v2

@@ 3 Characterisation analysis
3.1 Reverse I-V curves
We first measured reverse I-V curves using the Keithley 2450 sourcemeter both at room temperature (prior to the cool-down), and after the cool-down to the mixing chamber base temperature, $T_{\mathrm{CSNT}}$ = 9.4 $\pm$ 0.2 mK. The reverse I-V curve is used to characterise the relationship between the current flowing through the SiPM and the applied bias voltage, and can be used to determine its breakdown voltage, defined as the minimum bias voltage above which the SiPM can operate in Geiger mode and detect single photons. Our main motivation for acquiring reverse I-V curves was to ensure that the SiPM had survived the cool-down to base temperature before we attempted to acquire full characterisation data.
At room temperature, the dark count rate of the SiPM is large enough such that breakdown can be observed when the SiPM is in dark environment. At cryogenic temperatures however, an external light source is required to trigger breakdown. A signal pulse generator (1 kHz, 999 µs pulse width) and a current-limiting resistor were used to bias the LED whilst the I-V curve was acquired. Both the room temperature and mixing chamber base temperature reverse I-V curves were measured between 0 - 35 V, in 300 mV steps, with a 10 ms time delay between subsequent voltage steps during the sweep. The current limit on the Keithley 2450 was set to 60 µA to avoid damage to the SiPM. Figure 5 (left) shows a comparison of the reverse I-V curve taken at room temperature prior to cool-down, versus at the mixing chamber base temperature post cool-down. We clearly observe the SiPM breakdown both pre and post cool-down, demonstrating that the SiPM survived the cool-down to $T_{\mathrm{CSNT}}$ = 9.4 $\pm$ 0.2 mK. In addition, we observe a clear shift downwards in the breakdown voltage (i.e., the onset of self-sustaining avalanches) between the room temperature measurement and after cool-down, as expected with decreasing temperature.
We observed that after acquiring the reverse I-V curve, the temperature of the mixing chamber plate increased to $\sim$80 mK. This was measured using a Ruthenium Dioxide (RuO2) sensor mounted onto the mixing chamber plate, which measures the temperature accurately down to 20 mK; below 20 mK, the mixing chamber temperature is measured with the CSNT sensor. Given that the SiPM is thermally-coupled to the mixing chamber plate prior to commencing the I-V scans, it is likely that the temperature of the SiPM itself also increased during the sweeps. It is therefore very difficult to determine the breakdown voltage using the reverse I-V curve at a guaranteed temperature of 9.4 mK.
Furthermore, using follow up measurements of the SiPM submersed in a liquid helium (LHe) dewar at $T$ = 4 K, we observed that the reverse I-V curve shape is strongly dependent on the delay time used between subsequent voltage steps during the sweep. Figure 5 (right) shows reverse I-V curves taken in the LHe dewar using a delay time of 10 ms (as was used for the reverse I-V curves taken at the mixing chamber base temperature and at room temperature), versus 1 s. Figure 5 shows that at $T=4$ K, there is a clear difference between the reverse I-V curves using delay times of 10 ms and 1 s, both in the absolute current in the linear pre-breakdown regime, and in the apparent breakdown voltage.
We initially chose a 10 ms delay time between voltage steps as we did not want to inject unnecessary heat into the dilution refrigerator, however, we did not anticipate that there would be such an impact in the reverse I-V curve shape. This is likely driven by an increase in the SiPM quenching resistance at cryogenic temperatures, which lengthens the microcell recharge and thus requires a longer delay time during the sweep in order to get an accurate current measurement. We therefore cannot use the reverse I-V curves to get an accurate measurement of the breakdown voltage, which we also cannot guarantee to be the value of the breakdown at $T_{\mathrm{CSNT}}$ = 9.4 $\pm$ 0.2 mK due to the observed heating during the sweep. As such, we estimate the breakdown voltage using an alternative method, described in Section 3.3.
Figure 5: Left: Reverse I-V curves taken at room temperature prior to cool-down (green), and at the mixing chamber base temperature post cool-down (magenta). Right: Reverse I-V curves taken whilst the SiPM was submerged in a LHe dewar at $T$ = 4 K, acquired using a delay time of 10 ms (blue), and 1 s (orange) between subsequent voltage steps, demonstrating the impact that the delay time has on both the absolute current value in the linear pre-breakdown regime, and on apparent breakdown voltage.
@@ 3.2 Single photon response and gain
We first characterise the single photoelectron (SPE) response of the SiPM. To do this, we plot two histograms corresponding to the pulse amplitude and pulse area for each dataset. As an additional data cleaning cut for the SPE analysis, we only consider pulses where the time difference between the pulse in question and the preceding pulse is $\geq 2$ µs. This is to reject pulses with amplitudes less than the single photon pulse, that can occur from correlated noise or from noise fluctuations on the tails of the previous pulse.
@@
Figure 6 shows the pulse amplitude (left) and pulse area (right) distributions of the SiPM whilst operated at $V_{\mathrm{bias}}$ = 32 V. In both histograms, there are distinct peaks for successive numbers of PE. The first peak is the 1 PE signal, i.e., observed output pulse following the detection of a single photon. Due to intrinsic noise, a SiPM still generates output pulses in the absence of light. The 1 PE peak we observe in our distributions is due to a combination of thermally-induced dark count noise, and noise from correlated delayed avalanches occurring at times later than the SiPM recharge time. These noise sources are discussed in detail in Sections 3.4 and 3.6 respectively. The successive peaks corresponding to >>1 PE we observe in our distributions are a result of correlated noise from direct crosstalk, which manifest as superpositions of the 1 PE waveform. This noise source is discussed further in Section 3.5.
@@
The mean SPE response is extracted from a gaussian fit to the first peak in each spectrum. The SPE response at $V_{\mathrm{bias}}$ = 32 V is calculated to be $\mu_{\mathrm{SPE,amp}}$ = 3.65 mV $\pm$ 0.42 mV and $\mu_{\mathrm{SPE,area}}$ = 0.15 V $\pm$ 0.02 V in amplitude and area respectively. The mean SPE amplitude is smaller than the raw mean amplitude due to the moving-average filter which is applied to the data prior to running the pulse finding algorithm which reduces the height of the pulses; this can be seen in Figure 4.
@@
Figure 6: Pulse amplitude (left) and area (right) distributions of the NUV-HD-cryo SiPM operated at $V_{\mathrm{bias}}$ = 32 V. Gaussian functions are fit to the first 3 peaks in each histogram. The first peak in each histogram corresponds to the 1 PE peak.
@@
The separation between each pair of consecutive peaks in Figure 6 (right) is constant and is proportional to the gain, $G$, defined as the amount of charge generated in an avalanche in response to each detected photon. Each microcell in the SiPM produces a highly quantised amount of charge when a photon is absorbed in the active volume, and this response is extremely uniform across all microcells. This unique property of SiPMs results in excellent charge resolution, and allows for a very precise measurement of the gain. The gain is expected to increase proportionally with higher $\Delta V$.
To determine the separation, $S$, we use the results of the gaussian fits to the remaining peaks in the pulse area spectrum. We fit out the separation between peaks as a parameter in a linear polynomial fit to the fitted means of the 1, 2, and 3 PE peaks. The separation is calculated to be $S_{\mathrm{area}}$ = 0.149 $\pm$ 0.0003 V at $V_{\mathrm{bias}}$ = 32 V. $S_{\mathrm{area}}$ is multiplied by the oscilloscope sampling period to convert the separation into units of charge, which is then divided by the electron charge to calculate the gain:
@@
$G=\frac{S_{\mathrm{area}}[V]\cdot t_{\mathrm{s}}[s]}{R_{\mathrm{scope}}[\Omega]\cdot G_{\mathrm{board}}\cdot q_{\mathrm{e}}[C]},\qquad$ (3.1)
where $R_{\mathrm{scope}}=50\ \Omega$ is the input resistance of the oscilloscope, $G_{\mathrm{board}}=130$ V/V is the gain of the custom amplifier board from the design specification, $q_{\mathrm{e}}=1.6\times 10^{-19}$ C is the electron charge of an electron, and $t_{\mathrm{s}}$ = 4 ns is the sampling period.
Figure 7 shows the measured gain as a function of $V_{\mathrm{bias}}$ ($\Delta V$). As expected, we observe a proportional increase in the gain with $\Delta V$. The measured gain ranges between $7\times 10^{5}$ and $1.2\times 10^{6}$ for 32 V $V_{\mathrm{bias}}<$ 34 V. At 77 K, the gain is reported to be approximately $G=3.2\times 10^{6}$ for the NUV-HD-cryo technology [5], which is a factor of 3.7 higher compared to our value of $G=8.6\times 10^{5}$ for the same overvoltage value (7 V) based on an interpolation between our data points [^1] The overvoltage is defined in Section 3.3.. This difference could be driven by a “freeze-out” of free carriers at ultra-low temperatures, a mechanism which has previously been reported at LHe temperatures [6], however we do not explore this further in this manuscript.
Figure 7: Gain of NUV-HD-cryo SiPM as a function of $V_{\mathrm{bias}}$. Statistical error bars are shown, however are not necessarily visible as the errors are small. Blue dashed line is interpolated between the data points.
3.3 Breakdown voltage
To determine the breakdown voltage ($V_{\mathrm{bd}}$) of the SiPM at $T_{\mathrm{CSNT}}$ = 9.4 $\pm$ 0.2 mK, we linearly extrapolate the mean SPE amplitude $\mu_{\mathrm{SPE,amp}}$ down to zero. Since the pulse amplitude is proportional to the prompt avalanche current, it is sensitive to the onset of self-sustaining avalanches and therefore can be used to estimate the breakdown voltage.
We first identify candidate 1 PE signals using the pulse finder algorithm described in Section 2.2. Similarly to the SPE analysis described in the previous section, to ensure that we obtain a clean sample of 1 PE signals, i.e., excluding recharge-limited pulses with amplitudes << 1 PE or noise fluctuations on the tail, we require that the time between the pulse in question and the previous pulse $\geq 2$ µs. We then search for the maximum amplitude of the pulse in a second filtered version of the waveform, which uses a much shorter MA window of 8 ns (2 samples), within the ToT region identified by the pulse finder.
We do not use the 80 ns MA filtered pulse amplitude for this analysis since the filter reduces the 1 PE pulse amplitude; a longer MA window is intentionally chosen for the pulse finder to more effectively filter out baseline noise, however this does result in a decrease in the pulse height. Using this amplitude in the extrapolation would result in an overestimation of $V_{\mathrm{bd}}$. We therefore use the 8 ns MA filtered amplitude, which is a much closer proxy to the actual height of the 1 PE amplitude. We also repeat the analysis using the raw waveform amplitude, however this results in a large uncertainty on the estimated value of $V_{\mathrm{bd}}$ due to the presence of additional noise in the raw waveforms. We instead use this result as a cross-check with the result obtained from the 8 ns MA filtered amplitude.
Similarly to the SPE characterisation, we fit a gaussian to the 1 PE peak amplitude distribution in order to determine $\mu_{\mathrm{SPE,amp}}$ at each bias voltage. We fit these data points with a linear function and extrapolate down to the value of $V_{\mathrm{bias}}$ at which $\mu_{\mathrm{SPE,amp}}$ goes to zero. Figure 8 shows $\mu_{\mathrm{SPE,amp}}$ (using the 8 ns MA filtered amplitude) as a function of $V_{\mathrm{bias}}$, including the linear fit and uncertainty band; we determine a breakdown voltage of $V_{\mathrm{bd}}$ = 25.70 ± 0.04 V with this method. Repeating this analysis with the raw waveform amplitude yields a breakdown voltage of $V_{\mathrm{bd}}$ = 25.47 ± 0.58 V. Since our two measurements are consistent within uncertainties, we report the 8 ns filtered MA amplitude result. From hereafter, we report all characterisation measurements both in terms of the bias voltage and in the overvoltage, defined as $\Delta V=V_{\mathrm{bias}}-V_{\mathrm{bd}}$.
The breakdown voltage of the SiPM is expected to decrease with lower temperatures. At 77 K, the breakdown voltage of the NUV-HD-cryo technology is expected to be $V_{\mathrm{bd}}$ = 27.1 V [17]. As the selected photon sensors for the upcoming DarkSide-20k dark matter experiment, the collaboration has tested 359040 NUV-HD-cryo SiPMs at 77 K. The DarkSide-20k collaboration has determined the average breakdown voltage at 77 K to be $V_{\mathrm{bd}}$ = 27.19 $\pm$ 0.05 V [2], which is consistent with the expected value reported in [17]. The single device we are testing originates from an engineering run used for pre-production assembly, and is of the same type as the DarkSide-20k SiPMs. Therefore, our measurement of $V_{\mathrm{bd}}$ = 25.70 $\pm$ 0.04 V at $T_{\mathrm{CSNT}}$ = 9.4 $\pm$ 0.2 mK is consistent with the expected behaviour of $V_{\mathrm{bd}}$ to decrease with lower temperatures for these devices.
Figure 8: Mean SPE amplitude $\mu_{\mathrm{SPE,amp}}$ as a function of $V_{\mathrm{bias}}$, using the maximum amplitude obtained from the 8 ns MA filtered waveform. Linear fit and uncertainty band are also shown. The breakdown voltage extracted from extrapolating $\mu_{\mathrm{SPE,amp}}$ down to zero is $V_{\mathrm{bd}}$ = 25.70 ± 0.04 V.
3.4 Dark count rate
SiPMs can exhibit both uncorrelated and correlated noise contributions. The dark count rate (DCR) is the main source of uncorrelated noise. Each dark count is initiated by a thermally-generated charge carrier, which triggers an avalanche in the active region through the same process as a photon-generated charge carrier; thus DCR events are indistinguishable from single photon signals. The DCR of the SiPM is expected to increase with $\Delta V$ [17], since a higher $\Delta V$ increases the probability of an avalanche being triggered. The DCR is also expected to increase as a function of temperature. At higher temperatures, the DCR is dominated by field-enhanced thermal generation of charge carriers. However, below a certain temperature, the reduction of the DCR reaches a plateau, where the DCR becomes dominated by tunnelling effects that are weakly temperature dependent.
We characterise the DCR of the NUV-HD-cryo SiPM using the unshadowing method, based on using the time distribution between two consecutive pulses described in [12]. Given the incredibly low DCR of this technology, we need to consider the time differences of pulses across separate waveforms in addition to time differences within waveforms. Assuming a DCR of 5 mHz/mm2 at 77 K [17], this corresponds to a DCR of 0.5 Hz for a single SiPM. If we were to acquire 600,000 randomly-triggered waveforms, we would expect only 30 dark counts. Thus, to only consider time differences within a waveform would require extremely long waveforms which is not feasible from a data acquisition perspective. We make the assumption that we lose a negligible number of pulses between waveforms as a result of acquiring the data using the Tektronix FastFrame Memory scheme [26] described earlier in Section 2.2, which has a maximum dead time of 2.5 µs per waveform segment.
@@ 3.4 Dark count rate
Figure 9 (left) shows the measured time distribution of consecutive pulses whilst the SiPM is operated at $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 6.3 V). The dip in the $dt$ distribution at close to the waveform acquisition window, 100 µs, can be attributed to a loss in the pulse finder efficiency for pulses that occur near the end or beginning of waveform that may be partially cut-off, as well as from the dead time between waveforms itself, showing that the loss of pulses between waveforms is not necessarily negligible. We also observe that if we apply a stricter (larger) $dt$ cut between the primary pulse and the preceding pulse, the effect of the dip is slightly enhanced.
@@
Figure 9 (right) shows the unshadowed time distribution following the procedure in [12]. The unshadowed distribution is the average number of secondary pulses occurring between the bin limits of $t_{i}$ and $t_{i+1}$ defined in Figure 9 (left), normalised per unit time to determine the pulse rate. The pulse rate distribution contains two contributions: the correlated delayed pulse rate, and the DCR. At large times, the rate of correlated delayed pulses is expected to disappear, leaving only the DCR. To extract the DCR, we compute the weighted mean of the asymptotic rate, between 1 s <dt<<dt< 8 s. We calculate the DCR to be 4.046 $\pm$ 0.217 mHz/mm2 at $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 6.3 V), the value of which is indicated by the horizontal dashed line in Figure 9 to guide the eye. Figure 10 shows the measured DCR as a function of $V_{\mathrm{bias}}$ ($\Delta V$). As expected, the DCR increases as a function of $\Delta V$, increasing by 25% for 6.3 V $\Delta V<$ 8.3 V.
[…]
We note that the DCR values we calculate here are also consistent with the measured DCR value of <<5 mHz/mm2 at 77 K from [17], which could suggest that for temperatures below 77 K, the DCR is approximately constant. However, this is based on our measurement of a single device at $T_{\mathrm{CSNT}}$ = 9.4 $\pm$ 0.2 mK; measurements of several devices at more temperature points between 77 K and our base temperature would be required to confirm this.
@@
Figure 9: Left: Measured time distribution (dtdt) of consecutive pulses at $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 6.3 V). Right: Unshadowed time distribution, based on dtdt distribution in left-hand plot. Horizontal dashed line indicates calculated DCR of 4.046 $\pm$ 0.217 mHz/mm2.
@@
Figure 10: Measured DCR as a function of $V_{\mathrm{bias}}$ ($\Delta V$). The DCR increases by 25% between 6.3 V $\Delta V<$ 8.3 V. Red line is interpolated between the data points.